A function is defined by the polynomial 𝑓(π‘₯) = 3π‘₯4 βˆ’ 4π‘₯3 βˆ’ 12π‘₯2 + 8. Find and classify all the stationary points of f(x).


1
Expert's answer
2021-06-08T12:01:17-0400

Let us find and classify all the stationary points of the function 𝑓(π‘₯) = 3π‘₯4 βˆ’ 4π‘₯3 βˆ’ 12π‘₯2 + 8. The function is differetiable in all points of the real line. Let us find the points "x" for which "f'(x)=0." Since "f'(x)=12x^3-12x^2-24x," we conclude that "12x^3-12x^2-24x=0" implies "12x(x^2-x-2)=0", and hence "12x(x+1)(x-2)=0". It follows that "x_1=-1, \ x_2=0,\ x_3=2" are the stationary points of the function 𝑓(π‘₯) = 3π‘₯4 βˆ’ 4π‘₯3 βˆ’ 12π‘₯2 + 8. Taking into account that the function "f" is continuous and "f'(-2)=-96<0,\ f'(-0.5)=7.5>0,"

"f'(1)=-24<0,\ f'(3)=144>0," we conclude that the function is increasing on the intervals "(-1,0)" and "(2,+\infty)," and the function is deccreasing on the intervals "(-\infty,-1)" and "(0,2)."

Therefore, "x_1=-1,\ x_3=2" are the points of minimum, and "x_2=0" is the point of maximum.